OpenStudy (baby456):

hey please help foe easy medal +fan square root of 9/16 +5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\sqrt{\frac{9}{16}} +5\] like that?

OpenStudy (baby456):

yes

OpenStudy (anonymous):

what is \(\sqrt{9}\)?

OpenStudy (baby456):

3

OpenStudy (anonymous):

right !

OpenStudy (anonymous):

and what is \(\sqrt{16}\)?

OpenStudy (triciaal):



OpenStudy (baby456):

is it rounded to 3/4

OpenStudy (anonymous):

no dear it is not "rounded" it is exactly \(\frac{3}{4}\)

OpenStudy (triciaal):

@satellite73 not your fan use both values

OpenStudy (anonymous):

\[\huge \sqrt{\frac{9}{16}}=\frac{\sqrt{9}}{\sqrt{16}}=\frac{3}{4}\]

OpenStudy (baby456):

ok thanks

OpenStudy (anonymous):

then add 5 i guess

OpenStudy (anonymous):

what the monkey does "use both values" mean?

OpenStudy (triciaal):

when you take the square root -3/4 and +3/4

OpenStudy (triciaal):

negative * negative = positive and positive * positive = positive

OpenStudy (anonymous):

???

myininaya (myininaya):

sqrt(9/16) is the principal square root of 9/16 it is only the positive (or 0 if we have sqrt(0)) square root

OpenStudy (triciaal):